OpenStudy (animelover91804):

What value of h makes the statement true? (17 + h) + (–9) = 17 + (5 + (–9)) A. –21 B. –5 C. 5 D. 21

OpenStudy (alexandervonhumboldt2):

(17 + h) + (–9) = 17 + (5 + (–9)) 17+(h+(-9))=17+(5+(-9)) any ideas?

OpenStudy (animelover91804):

I gots it